OpenStudy (anonymous):

Solve the following system using the elongation method. 2a+2b=4 4a-3b=8

RhondaSommer (rhondasommer):

You mean elimination?

OpenStudy (anonymous):

Yes sorry!!

RhondaSommer (rhondasommer):

you would multiple the top equation \[(2a+2b=4)\] by negative 2 to get \[ -4a-4b=-8\] then you combine them \[-4a-4b=-8\]\[4a-3b=8\] the -4a and 4a eliminate. however the -8 and 8 eliminate as well so this equation would be \[-7b=0\]so \[b=0\] :)

OpenStudy (anonymous):

Thank you so much!!

RhondaSommer (rhondasommer):

then, you place it into the origional equation. 2a+2(0)=4 2a=4 a=2 :)